OpenStudy (anonymous):

Which is a factor of 28xy – 154x + 20y – 110 ?

OpenStudy (callisto):

28xy – 154x + 20y – 110 = 14x ( 2y - 11) + 10 ( 2y - 11) = (14x + 10)(2y - 11) = 2 ( 7x + 5)(2y - 11)

OpenStudy (anonymous):

thanks love 

OpenStudy (callisto):

Welcome. Basically, you just take out the common factor, group the terms and repeat these steps until you get the factored expression.

OpenStudy (anonymous):

alright thanks hun if i have any ther questions ill ask :)